# Measurement induced phase transition in a quantum Ising model by periodic measurements

### v1

@@ Abstract
Measurement-induced phase transitions are often studied in random quantum circuits, with local measurements performed with a certain probability. We present here a model where a global measurement is performed with certainty at every time-step of the measurement protocol. Each time step, therefore, consists of evolution under the transverse Ising Hamiltonian for a time $\tau$, followed by a measurement that provides a “yes/no” answer to the question, “Are all spins up?”. The survival probability after $n$ time-steps is defined as the probability that the answer is “no” in all the $n$ time-steps. For various $\tau$ values, we compute the survival probability, entanglement in bipartition, and the generalized geometric measure, a genuine multiparty entanglement, for a chain of size $L\sim 26$, and identify a transition at $\tau_{c}\sim 0.2$ for field strength $h=1/2$. We then analytically derive a recursion relation that enables us to calculate the survival probability for system sizes up to 1000, which provides evidence of a scaling $\tau_{c}\sim 1/\sqrt{L}$. The transition at finite $\tau_{c}$ for $L\sim 28$ seems therefore to recede to $\tau_{c}=0$ in the thermodynamic limit. Additionally, at large time-steps, survival probability decays logarithmically only when the ground state of the Hamiltonian is paramagnetic. Such decay is not present when the ground state is ferromagnetic.
I Introduction
@@
A paradigmatic model for studying measurement-induced phase transition (MIPT) is a quantum gate circuit of some suitable structure in which measurements are performed locally Li et al. (2018); Skinner et al. (2019); Rakovszky et al. (2019); Li et al. (2019); Chan et al. (2019); Szyniszewski et al. (2019); Bao et al. (2020); Choi et al. (2020); Gullans and Huse (2020a, b); Jian et al. (2020); Zabalo et al. (2020); Iaconis et al. (2020); Turkeshi et al. (2020); Zhang et al. (2020); Szyniszewski et al. (2020); Nahum et al. (2021); Ippoliti et al. (2021); Ippoliti and Khemani (2021); Lavasani et al. (2021a, b); Sang and Hsieh (2021); Block et al. (2022); Sharma et al. (2022); Agrawal et al. (2022); Barratt et al. (2022); Jian et al. (2023); Kelly et al. (2023), (for review, see Fisher et al. (2023)). The time evolution of the system consists of unitary evolution under some operator $U$ for an interval of time, followed by (projective or weak) measurement of some suitable local operator with probability ${\mathcal{P}}$. After every time step, the bipartite entanglement entropy $S(t)$ is measured. It is found that in the asymptotic limit of large size and long time, for small values of ${\mathcal{P}}$, $S(t)$ increases linearly with time, while for large values of ${\mathcal{P}}$ it saturates after some time. At some intermediate value ${\mathcal{P}}={\mathcal{P}}_{c}$, the entropy increases logarithmically with time, and this is considered to be a signature of a phase transition that has been induced by the process of measurement. The occurrence of such transitions has been confirmed in numerous systems in the last 5 years, and most of them are ‘small’ in size ($<40$) Skinner et al. (2019); Chan et al. (2019); Szyniszewski et al. (2019); Bao et al. (2020); Choi et al. (2020), while some are on larger sizes ($>100$) Li et al. (2018, 2019); Gullans and Huse (2020a, b). The location of the transition point ${\mathcal{P}}_{c}$ is believed to be independent of size. Apart from the random circuit analysis, this type of transition is also captured in the transverse field Ising model  Boorman et al. (2022); Sierant et al. (2022a); Malakar et al. (2024); Su et al. (2024); Tirrito et al. (2023); Roser et al. (2023); Biella and Schiró (2021); Banerjee and Sengupta (2024); Paviglianiti and Silva (2023), fermionic systems Buchhold et al. (2021); Poboiko et al. (2024); Müller et al. (2022); Chatterjee and Modak (2024); Minato et al. (2022); Jin and Martin (2024); Jian et al. (2021), many-body localized systems Lunt and Pal (2020) and higher-dimensional systems Turkeshi et al. (2020); Lunt et al. (2021); Nahum et al. (2021); Sierant et al. (2022b); Feng et al. (2023). This volume-to-area law transition is also observed in experiments with trapped ions and superconducting qubits Noel et al. (2022); Koh et al. (2023); AI and Collaborators (2023); Kamakari et al. (2024). Dependence of MIPT on the entangling power of unitary gate has been explored in Manna et al. (2024). Two essential ingredients of these studies are the local nature of the measurement and the randomness in whether the measurement will be performed or not at a particular time step. It would therefore be interesting to investigate what happens when these two features are removed. More importantly, for an integrable system, if one can go to a larger size ($\sim 1000$), it would be useful to check if the transition point depends on system size.
@@
In this work, we shall consider non-random global projective measurement performed on a quantum Ising chain evolving under the transverse Ising Hamiltonian. We start from a chosen pure state $|\psi_{0}\rangle$ and one time step consists of evolution for time $\tau$ followed by the measurement. (Note that the measurement is performed with certainty at every time step, and the only source of stochasticity is the quantum nature of the system.) We study, as a function of time, three quantities, namely, bipartite entanglement, multipartite entanglement, and the survival probability of the initial state $|\psi_{0}\rangle$. From numerical studies on chains of size $L\leq 28$, it is found that (i) the bipartite entanglement shows a volume-to-area law transition at some critical value $\tau=\tau_{c}$ and (ii) the multipartite entanglement as well as the survival probability also show non-analytic behavior at the same value of $\tau$. Thus, we observe a ‘measurement-induced phase transition’ just like random circuit models Skinner et al. (2019); Koh et al. (2023). This transition point exists for every value of $h$ and is hence unrelated to the order-disorder transition at $h=1$. We present below the results for $h=\frac{1}{2}$ and include the results for $h=\frac{3}{2}$ in the appendix A. The central result of this work is that, exploiting the analytic nature of the Hamiltonian, we can calculate the survival probability at large $L$($\sim 1000$) too and show that the critical value $\tau_{c}$ scales with system size as $1/\sqrt{L}$. This proves that $\tau_{c}$ vanishes at the thermodynamic limit, and the transition we observe at small size exists for finite sizes only. Hence, whether the transition in random circuit models survives in the thermodynamic limit is an important question.
In the next section II, we present the measurement protocol. In Sections III and IV we present the analysis of bipartite entanglement and generalized geometric measure in small systems. Sec. V is devoted to the analytic theory, and Sec. VI is devoted to the summary.
II protocol and survival probability
Figure 1: $R_{n}$ vs $n$: Survival probability for different $\tau$ values : Here, we have plotted $R_{n}$ vs. $n$ graph corresponding to each $\tau$ value for system size $L=28$ at field $h=\frac{1}{2}$ where $R_{n}$ is the survival probability (Eq. 7) and $n$ is the number of cycle of the protocol.
@@
It is allowed to evolve unitarily for a time $\tau$ under the transverse Ising Hamiltonian.
$\mathcal{H}=-J\sum_{j=1}^{L}s^{x}_{j}s^{x}_{j+1}-\Gamma\sum_{j=1}^{L}s^{z}_{j}$ (2)
(where $J$ is the interaction strength and $\Gamma$ is the strength of the transverse field. We fix $J=1$ and use symbol $h$ for $\Gamma/J$ ).
We now perform a projective measurement of the transverse magnetic moment per site $M_{z}=(1/L)\sum_{i=1}^{L}s_{i}^{z}$ and ask the question
Figure 2: Behavior of height of the plateau region of $R_{n}$ with $\tau$: In panel (a), we have plotted the $H$ vs. $\tau$ graph, where $H$ is denoted as the height of the plateau region for a given $\tau$ that is already depicted in FIG. 1. When there are multiple plateaus, we consider the height of the first one. Here, the black color denotes the curve corresponding to field $h=0.5$, and the orange color denotes the curve corresponding to $h=\frac{3}{2}$. The transition occurs around $\tau_{c}=0.2$. The point where this transition occurs is perfectly realized from the curve of panel (b), where we have plotted $\frac{dH}{d\tau}$ vs. $\tau$ for various system sizes, that is, $L=20,22,24,26$ at field $h=\frac{1}{2}$. Here, the highest value to each curve corresponds to the transition point.
Is the magnetic moment (per site) $M_{z}=1$? (3)
We perform the experiment on a large number ${\mathcal{N}}$ of identically prepared replicas, then note down the fraction of cases where the answer is yes and call this fraction $p_{1}$. We discard the replicas that gave answer yes and with the other replicas carry on the unitary evolution under ${\mathcal{H}}$ for another time interval $\tau$ (in units of $J/\hbar$). Then we perform the same measurement and note down the number of replicas, say, $p_{2}{\mathcal{N}}$, which yield yes. We repeat this procedure of unitary evolution, followed by projective measurement, so that after $n$ such cycles we get the number of replicas, say, $p_{n}{\mathcal{N}}$, which yield yes. Then, the probability that the answer is yes for the first time at the $n$-th measurement is $p_{n}$, and the probability that the system has yielded no at the first $n$ successive measurements is
@@
We shall call $R_{n}$ the survival probability. The above protocol for dynamics has been used earlier Dhar et al. (2015a, b); Dhar and Dasgupta (2016) as a reasonable protocol for measuring the probability of first occurrence within the framework of quantum measurement.
Figure 3: $\tau_{c}$ vs $L$ and $h$ : Here, the critical point $\tau_{c}$ as we obtain from the survival probability $R_{n}$ is plotted with respect to the system size $L$ and field strength $h$.
Analysis for small systems: After the time evolution from $t=0$ to $t=\tau$, the system attains a state $\exp(-i\mathcal{H}\tau)|\psi_{0}\rangle$. Since $|\psi_{0}\rangle$ is the only state with $M_{z}=1$, the effect of the measurement stated in (3) is to project out $|\psi_{0}\rangle$ and get the state
@@
The state function $|\psi_{n}\rangle$ after $n$ states is calculated numerically using the Chebyshev polynomial technique Piotr et al. (2019). and the survival probability is obtained therefrom as
$R_{n}=\langle\psi_{n}|\psi_{n}\rangle$ (7)
The plot of $R_{n}$ as a function of $n$ (for a given $\tau$) shows a plateau region at small $n$. As we increase the value of $\tau$, this region increases in value and saturates near 1 for a high value of $\tau$ (FIG. 1).
The principal quantity of our study is the rate at which the height of the plateau region varies with $\tau$ (FIG. 2(a)), and we observe that this quantity shows a peak at some value of $\tau$ (FIG. 2(b)).
Figure 4: Survival probability $R_{n}$ vs time-step $n$ : Here, the survival probability $R_{n}$ is plotted with respect to time-step $n$ for $h=\frac{3}{2}$. System size is $L=50$. A logarithmic decay at large time-steps is observed, which is not present at $h<1$.
When there are multiple plateaus, we consider the height of the first one. We note that the height of the peak rises with system size and identify this value of $\tau$ as the critical value $\tau_{c}$ (FIG. 2(b)) where the plot of $\frac{dH}{d\tau}$ vs $\tau$ is given. Throughout the paper, all the derivatives are calculated numerically by the two-point formula. We note that at finite size we have different $\tau_{c}$ at $h=\frac{1}{2}$ and $h=\frac{3}{2}$. The dependence of critical point $\tau_{c}$ on system size and $h$ is studied in FIG (3). It is observed that the dependence of $\tau_{c}$ on $h$ decreases as the system size $L$ increases. $\tau_{c}$ also moves towards smaller values as system size $L$ increases.
In FIG (4), the survival probability is studied at large $n$ at $h=\frac{3}{2}$ for different $\tau$ values. We can see that $R_{n}$ decays in logarithmic fashion at large $n$. This kind of logarithmic decay is not present in ferromagnetic phase ($h<1$) and can be observed only when the system is in paramagnetic phase ($h>1$). However, when $h$ is too high, such decay is not present. Thus, the pattern of decay can detect the different quantum phases of the ground state of transverse field Ising model.
III bipartite entanglement
Figure 5: Entanglement behavior with system size $L$ and $\tau$ : In panel (a), entanglement entropy for the whole system in the $L/4$ : rest bipartition ($S(L/4)$) is plotted for different $\tau$ values with respect to $L$, which is the system size. $\tau$ values increase as the color of the curves becomes deeper. We have taken $n=10$ for measuring $S$. It is apparent that there is an area law to volume law transition. In panel (b) we have plotted $S(L/4,\tau)$ for different values of $L$ with respect to $\tau$. In panel (c) we have shown the scaling collapse of the entanglement curves occurs when $\tau_{c}=0.2$ with appropriate exponents, which is $0.3$ following Ref. Skinner et al. (2019); Koh et al. (2023).
Bipartite entanglement refers to entanglement, a quantum correlation Horodecki et al. (2009); Gühne and Tóth (2009); Das et al. (2017) between two subsystems (or parties) in a composite quantum system. Suppose that $A$ and $B$ are two parts of a single quantum system, and that the state of the entire system is $\rho^{AB}$.The state is said to be entangled if it cannot be decomposed into $\rho^{AB}=\sum_{i}p_{i}\rho^{A}_{i}\otimes\rho^{B}_{i}$, for any ensemble $\{p_{i},\rho^{A}_{i}\otimes\rho^{B}_{i}\}$. For calculating the bipartite entanglement of a pure bipartite state, the measure we consider is the von Neumann entropy or entanglement entropy Bennett et al. (1996). Let us consider the two subsystems. $A$ and $B$ consist of $\ell$ and $L-\ell$ sites, respectively. Here, $L$ is the total number of sites. The measure, entanglement entropy, for any pure state of the Ising chain, after $n$ time-steps, is computed as
@@
where $\rho_{A}$ and $\rho_{B}$ are the reduced density matrices for the subsystems $A$ and $B$. Throughout the paper, we have taken log base e when we calculate bipartite entanglement. In this logarithmic base, the entanglement of bell states that is maximally entangled states is $0.693$ nats.
Following Srednicki (1993); Eisert et al. (2010), we know that the entanglement entropy of non-degenerate ground state of a gapped Hamiltonian, follow the area law Hastings (2007) (i.e., it saturates as the subsystem size increases), while a generic state of the system follows the volume law (i.e. changes linearly with subsystem size).
In the context of MIPT, a volume law to area law transition of the entanglement entropy occurs at a critical rate of measurement ($\mathcal{P}_{c}$) as defined in Ref. Skinner et al. (2019). According to Skinner et al. (2019), near the critical point, for certain set-ups, there exists a general scaling entanglement entropy, which can be written as $S(\mathcal{P})-S(\mathcal{P}_{c})=F\left((\mathcal{P}-\mathcal{P}_{c})L^{\frac{1}{\nu}}\right)$ where $L$ is the system size and $S$ is the entanglement entropy of the whole system in the $L/2:L/2$ bipartition, which is the von Neumann entropy of a subsystem of size $L/2$. The form of $F$ would depend on the set-up considered. The entanglement entropy depends more strongly on system size in the volume law phase, than in the area law phase. When $S(\mathcal{P})-S(\mathcal{P}_{c})$ is plotted against $(\mathcal{P}-\mathcal{P}_{c})L^{1/\nu}$, all the curves for different $L$ collapse on each other around $\mathcal{P}_{c}$, which indicates a phase transition at $\mathcal{P}_{c}$.
In this section, we have calculated the entanglement in the $L/4:\text{rest}$ biseparation of the entire system, and referred to it as $S(L/4,\tau)$, for different $\tau$ values Li et al. (2018). The entanglement is quantified by using the entanglement entropy, i.e., the von Neumann entropy of the $L/4$ subsystems, for different $L$. An area law to volume law transition can be observed when $S(L/4,\tau)$ is plotted with respect to $\tau$ (FIG. 5(a)). In FIG. 5(b), entanglements for different $L$ collapses to similar values at low $\tau$, and becomes system-size dependent at larger $\tau$, which also indicates an area law to volume law transition. In FIG. 5(c), $S(L/4,\tau)-S(L/4,\tau_{c})$ is plotted with respect to $(\tau-\tau_{c})L^{\mu}$. We find that scaling collapse occurs around $\tau_{c}=0.2$ and $\mu=0.3$, which shows that an entanglement area law phase transition point exists around $\tau_{c}=0.2$ (FIG. 5).
IV GENERALIZED GEOMETRIC MEASURE
For quantifying multipartite entanglement, we utilize the measure referred to as the generalized geometric entanglement (GGM) Sen (De); Das et al. (2016). See also Barnum and Linden (2001); Plenio and V. (2001); Wei and Goldbart (2003); Osterloh and Siewert (2005); Yu and Song (2006); Orús (2008). A pure multiparty quantum state is said to be genuinely multiparty entangled if it cannot be factorized into a product state in any biseparation. The GGM of any pure multiparty state is defined as the minimum distance of the given state from the set of all states that are not genuinely multiparty entangled, with the Bures metric being used to obtain the distance function. We are interested in analyzing the genuine multiparty entanglement of the pure $L$-partite state, $|\psi_{n}\rangle$, after the $n$th measurement.
The GGM of |$\ket{\psi_{n}}$ is defined as
$G(n)=1-\Lambda^{2}_{max}(\ket{\psi_{n}}),$ (9)
where $\Lambda_{max}(\ket{\psi_{n}})$ is the maximum fidelity between the given state |$\ket{\psi_{n}}$ and an arbitrary non-genuinely multiparty entangled state |$\ket{\xi}$. So, a maximization is to be performed over all non-genuinely multiparty entangled pure states, which, however, can be explicitly obtained in closed form Sen (De):
$G(n)=1-\max_{\ell}\left\{\lambda_{\ell,L-\ell}^{2}\right\}_{l}$ (10)
where $\lambda_{\ell,L-\ell}$ is the maximal Schmidt coefficient in the bipartite split $\ell:L-\ell$ of |$\ket{\psi_{n}}$.
Since we calculate the GGM for the state $|\psi_{n}\rangle$ after the $n$th measurement, we plot the “stochastically attained GGM” (SAG), defined as
$S_{G}(n)=R_{n}\times G(n)$
The plot of $S_{G}(n)$ with $n$ at field strength $h=1/2$ is depicted in FIG. 6(a) for different $\tau$ values, viz. $\tau=0.1,0.2,0.3,\ldots,1.0.$ A similar trend, that is, a large gap between the curves at $\tau$ values $0.1$ and $0.2$, is noticed, like the $R_{n}$ vs $n$ graph in FIG. 1. To further analyze the behavior of multipartite entanglement, we perform cumulative averaging, which, for a sequence, $\bigl{\{}x_{1},x_{2},x_{3},\ldots,x_{n}\bigl{\}}$, is defined as $\mathcal{M}_{i}=\frac{1}{i}\sum_{k=1}^{i}x_{k}$. Therefore, $\mathcal{M}_{i}$ is the cumulative average at the $i$th step.
In our case, the cumulative average of $S_{G}(n)$ is denoted as $\mathcal{G}_{n}$, and in FIG. 6(b), we have plotted the $\mathcal{G}_{n}$ against $n$ for the same set of $\tau$ values as in FIG. 6(a) for the same field strength, that is $h=\frac{1}{2}$. The gap between $\tau=0.1$ and $\tau=0.2$ curves becomes more prominent after we consider the cumulative values. For $\tau=0.2$ and onwards, the gap starts to decrease, and from $\tau=0.3$, all the curves overlap on each other. We point out that the curves show a plateau region, just like survival probability for small $n$ (FIG. 1). This feature of the curves, corresponding to different values of $\tau$, leads us to check the behavior of $\frac{d\mathcal{G}}{d\tau}$ with $\tau$. We identify the value of $\tau$ corresponding to the highest value of the derivative $\frac{d\mathcal{G}}{d\tau}$ as $\tau_{c}$.
Thus, for sizes $L\leq 28$, the three quantities, namely survival probability, bipartite entanglement entropy, and genuine multipartite entanglement, show a critical behavior at a common value of $\tau_{c}$ around $0.2$ at field strength $h=1/2$, and this we consider as the signature of a measurement-induced transition in this quantum many-body system. These results resemble those in the studies on random circuit systems with local stochastic measurements Koh et al. (2023); Skinner et al. (2019). We have repeated the entire analysis for all the three mentioned order parameters at field strength, $h=\frac{3}{2}$, and the results are elaborated in Appendix A.

### v3

# Disappearance of measurement-induced phase transition in a quantum spin system for large sizes

@@ Abstract
Quantum measurements cause an abrupt change, called measurement induced phase transition, in the scaling behaviour of entanglement entropy in a quantum many-body system. The phenomenon is often studied in random quantum circuits, with local measurements performed with a certain probability. We report here one such transition in an interacting spin 1/2 system where a global measurement is performed with certainty at every time-step. We start with a pure state with all the spins up. Each time step consists of evolution under the transverse Ising Hamiltonian for a time $\tau$, followed by a measurement that provides a “yes/no” answer to the question, “Are all spins up?”. For various $\tau$ values, we compute the survival probability of the inital state, entanglement in bipartition, and the generalized geometric measure (a genuine multiparty entanglement), for a chain of size $L\sim 28$, and identify the transition points $\tau_{c}$ for different field strengths. We then analytically derive a recursion relation that enables us to calculate the survival probability for system sizes up to 1000, and provides evidence of a scaling $\tau_{c}\sim 1/\sqrt{L}$. Therefore, the transition at finite $\tau_{c}$ for $L\sim 28$ seems to recede to $\tau_{c}=0$ in the thermodynamic limit. Our work prompts a widespread investigation on the existence of measurement-induced transitions at large system-size.
@@
Measurement hypothesis of Quantum Mechanics indicates that measurement of a variable on a quantum state causes the state to collapse to an eigenstate of the measurement operator. When this change is studied as a function of the control parameter, sometimes it so happens that the change is qualitatively different on two sides of some special values of the parameters. Such phenomena are called measurement-induced phase transition, in analogy with the change in physical properties of a thermodynamic system at a phase transition. A paradigmatic model for studying measurement-induced phase transition (MIPT) is a quantum gate circuit of some suitable structure in which measurements are performed locally Li et al. (2018); Skinner et al. (2019); Rakovszky et al. (2019); Li et al. (2019); Chan et al. (2019); Szyniszewski et al. (2019); Bao et al. (2020); Choi et al. (2020); Gullans and Huse (2020a, b); Jian et al. (2020); Zabalo et al. (2020); Iaconis et al. (2020); Turkeshi et al. (2020); Zhang et al. (2020); Szyniszewski et al. (2020); Nahum et al. (2021); Ippoliti et al. (2021); Ippoliti and Khemani (2021); Lavasani et al. (2021a, b); Sang and Hsieh (2021); Block et al. (2022); Sharma et al. (2022); Agrawal et al. (2022); Barratt et al. (2022); Jian et al. (2023); Kelly et al. (2023), (for review, see Fisher et al. (2023)). The time evolution of the system consists of unitary evolution under some operator $U$ for an interval of time, followed by (projective or weak) measurement of some suitable local operator with probability $\mathcal{P}$}. After every time step, the bipartite entanglement entropy $S(t)$ is measured. It is found that in the asymptotic limit of large size and long time, for small values of $\mathcal{P}$}, $S(t)$ increases linearly with time (“volume law”), while for large values of $\mathcal{P}$} it saturates after some time (“area law”). At some intermediate value ${\mathcal{P}}={\mathcal{P}}_{c}$, the entropy increases logarithmically with time, and this is considered to be a signature of a phase transition that has been induced by the process of measurement. The occurrence of such transitions has been confirmed in numerous systems in the last five years, and most of them are ‘small’ in size (<40<40) Skinner et al. (2019); Chan et al. (2019); Szyniszewski et al. (2019); Bao et al. (2020); Choi et al. (2020); Manna et al. (2024), while some are larger (>100>100) Li et al. (2018, 2019); Gullans and Huse (2020a, b). The location of the transition point ${\mathcal{P}}_{c}$ is believed to be independent of size. Apart from the random circuit analysis, this type of transition is also captured in the transverse field Ising model  Boorman et al. (2022); Sierant et al. (2022a); Malakar et al. (2024); Su et al. (2024); Tirrito et al. (2023); Roser et al. (2023); Biella and Schiró (2021); Banerjee and Sengupta (2024); Paviglianiti and Silva (2023), fermionic systems Buchhold et al. (2021); Poboiko et al. (2024); Müller et al. (2022); Chatterjee and Modak (2024); Minato et al. (2022); Jin and Martin (2024); Jian et al. (2021); Han-Ze et al. (2025), many-body localized systems Lunt and Pal (2020), higher-dimensional systems Turkeshi et al. (2020); Lunt et al. (2021); Nahum et al. (2021); Sierant et al. (2022b); Feng et al. (2023) and in experiments with trapped ions and superconducting qubits Noel et al. (2022); Koh et al. (2023); AI and Collaborators (2023); Kamakari et al. (2024). This volume-to-area law transition basically arises due to the competition between the unitary evolution, which increases the entanglement, and the process of measurement, which decreases the entanglement. At the same time, ref. Coppola et al. (2022) shows a measurement-induced phase transition in a one-dimensional quadratic fermionic model under local and probabilistically applied measurement in finite system size and shows in the thermodynamic limit this transition vanishes and only the area law phase persists. Absence of MIPT has also been shown for random projective measurements Poboiko et al. (2023), for open 1D fermion system Starchl et al. (2025), for Dirac fermions Müller et al. (2025) and in dissipative quantum Ising chain Piccitto et al. (2022).
@@
In this work, we shall consider non-random global projective measurement performed on a quantum Ising chain evolving under the transverse Ising Hamiltonian. We start from a chosen pure state $|\psi_{0}\rangle$ and one time step consists of evolution for time $\tau$ followed by a measurement. (Note that the measurement is performed with certainty at every time step, and the only source of stochasticity is the quantum nature of the system.) At each time-step, we compute three quantities, namely, the bipartite entanglement, the multipartite entanglement, and the survival probability of the initial state $|\psi_{0}\rangle$. From numerical studies on chains of size $L\leq 28$, it is found that (i) the bipartite entanglement shows a volume-to-area law transition at some critical value $\tau=\tau_{c}$ and (ii) the multipartite entanglement as well as the survival probability also bear the signature of a transition at the same value of $\tau$. Thus, we observe a ‘measurement-induced phase transition’ just as in the case of random circuit models Skinner et al. (2019); Koh et al. (2023). The central result of this work is that, exploiting the analytic nature of the Hamiltonian, we can calculate the survival probability at large $L$ ($\sim 1000$) too and show analytically (and numerically) that the critical value $\tau_{c}$ scales with system size as $1/\sqrt{L}$. This proves that $\tau_{c}$ vanishes at the thermodynamic limit, and that the transition we observe at small size exists for finite sizes only. Hence, whether the transition in random circuit or other similar models survive in the thermodynamic limit is an important question.
Incidentally, we also find that in contrast to the random circuit models, our measurement does not always reduce the entanglement, and the transition occurs despite this fact. Furthermore, the transition point seems to exist irrespective of the presence of order in the ground state of the Hamiltonian; we present here the results for $h=1/2$ in the main paper (where the ground state is ordered) and include the (qualitatively similar) results for $h=3/2$ (where the ground state is disordered) and $h=1$, (equilibrium transition point) in the Supplementary Material I. Therefore, the equilibrium critical point has no effect on our observation. We believe that our protocol can be implemented in such experimental setups, where global measurement is performed in multiparty entangled systems. One such experimental work is demonstrated in Ref. Alexander and Mahesh (2025), where global measurement is performed in a highly entangled system.
We shall first present the measurement protocol and then the numerical results for the survival probability, bipartite entanglement and generalised geometric measure. Then we shall present the analytic results, followed by summary.
Figure 1: (a) Survival probability $R_{n}$ (as defined in Eq. 3) as a function of time-step $n$ for different $\tau$ values at $L=28$ and $h=\frac{1}{2}$. In the main figure, we consider $\tau$ values ranging from $\tau=0.1$ to $\tau=0.5$, taken at intervals of $0.1$. The inset shows same $R_{n}$ vs. $n$ plots corresponding to system size $L=28$ for $\tau$ values ranging from $\tau=0.1$ to $\tau=1.0$, in intervals of of $0.1$. The color bar represents the same $\tau$ values in both the main figure and the inset:(b) shows $dH/d\tau$ vs, $\tau$ plot for time-step $n=10$. The derivative $dH/d\tau$ shows a peak at $\tau=\tau_{c}$ the location of which varies slightly with system-size and is around 0.2 for $L=28$.
Protocol and Survival Probability:
@@
It is allowed to evolve unitarily for a time $\tau$ under the transverse Ising Hamiltonian
$\mathcal{H}=-\sum_{j=1}^{L}s^{x}_{j}s^{x}_{j+1}-h\sum_{j=1}^{L}s^{z}_{j}$ (2)
We now perform a projective measurement which provides a binary (yes/no) answer to the question “Are all spins up?”.
We perform the experiment on a large number $\mathcal{N}$} of identically prepared replicas, then note down the fraction of cases where the answer is yes and call this fraction $p_{1}$. We discard the replicas that gave answer yes and with the other replicas carry on the unitary evolution under $\mathcal{H}$} for another time interval $\tau$. Then we perform the same measurement and note down the number of replicas, say, $p_{2}{\mathcal{N}}$, which yield yes. We repeat this procedure of unitary evolution, followed by projective measurement, so that after the $n$-th cycle, we get the number of replicas, say, $p_{n}{\mathcal{N}}$, which yield yes. Then the probability that the system has yielded no for the first $n$ successive measurements is
@@
We shall call $R_{n}$ the survival probability as the initial state has escaped detection over $n$ steps with this probability. The above protocol for dynamics has been used earlier Dhar et al. (2015a, b); Dhar and Dasgupta (2016) as a reasonable protocol for measuring the probability of first occurrence within the framework of quantum measurement.
After the time evolution from $t=0$ to $t=\tau$, the system attains a state $\exp(-i\mathcal{H}\tau)|\psi_{0}\rangle$ and the effect of the measurement is to project out $|\psi_{0}\rangle$ and get the state
@@
Following this principle, the state function $|\psi_{n}\rangle$ after $n$ steps can be calculated numerically. We have used the Chebyshev polynomial technique Piotr et al. (2019) to facilitate the computation and studied entanglement and survival probability for the state $|\psi_{n}\rangle$.
The plot of $R_{n}$ as a function of $n$ (for a given $\tau$) shows a plateau region at small $n$. As we increase the value of $\tau$, the height of this plateau approaches unity (Fig. 1(a)). The principal quantity of our study is the rate at which the height of the plateau region varies with $\tau$ (Fig. 1(a)). The precise definition of this height is as follows. The $R_{n}$ vs $n$ curve shows at least one range of values of $n$ (say $n_{1}<n<n_{2}$) where the plot is horizontal and $R_{n}$ does not vary with $n$. The value of $R_{n}$ in this region is identified as the height of the plateau. In some cases there are several such regions and then we choose the first region (that is, the one for which $n_{1}$ and $n_{2}$ are smallest) for measuring the height of the plateau. We describe the height (H) of the plateau pictorially in Fig. 1(a). We observe that the derivative of this height with respect to $\tau$ shows a peak at some value of $\tau$ (Fig. 1(b)). Thus, for equispaced values of $\tau$, the vertical separation of the plateau regions is larger for intermediate values of $\tau$. We note that the height of the peak rises with system size and identify this value of $\tau$ as the critical value $\tau_{c}$ (Fig. 1(b)). We shall now see that at $\tau_{c}$ the entanglement entropy also shows a typical transition.
Bipartite Entanglement:
The system of $L$ sites is divided into two subsystems A and B, consisting of $\ell$ and $L$-$\ell$ (consecutive) sites. The entanglement entropy for a pure state is computed as
@@
where $\rho_{A}$ and $\rho_{B}$ are the reduced density matrices for the subsystems $A$ and $B$. In the context of MIPT, a volume law to area law transition of the entanglement entropy at some critical rate of measurement ($\mathcal{P}_{c}$) is identified Skinner et al. (2019) by using a scaling relation
$S(\mathcal{P})-S(\mathcal{P}_{c})=F\left((\mathcal{P}-\mathcal{P}_{c})L^{1/\nu}\right)$
Choosing $\ell=L/4$, and using a scaling function
$S(\tau)-S(\tau_{c})=F\left((\tau-\tau_{c})L^{{1/\nu}}\right)$ (7)
In Fig. 2 (a)), considering various $\tau$ values, we observe that the curves remain nearly flat for $\tau<0.20$, while for $\tau>0.20$ a linear growth of entanglement emerges. This behavior clearly indicates a transition from an area-law to a volume-law scaling, occurring around $\tau_{c}\approx 0.20$}. we also find a collapse (Fig. 2 (b)) with $\tau_{c}=0.2$ and $\nu=1.013$, thus, confirm the presence of an area-to-volume law transition more robustly. A similar study based on bipartite entanglement entropy for $h=\frac{3}{2}$ and h=1h={1} is provided in the supplementary material I, an area-to-volume law transition occurs about $\tau_{c}=0.1$ and $\tau_{c}=0.16$ for $h=\frac{3}{2}$ and h=1h={1} respectively. An additional confirmation for the transition is observed from the collapse of the plot of $H(\tau)-H(\tau_{c})$ vs $(\tau-\tau_{c})L^{1/\lambda}$ (Fig. 3), where $\lambda=0.4317$.
Figure 2: Behaviour of the bipartite entanglement entropy with system size $L$ and $\tau$ for $h=1/2$. In panel (a), entanglement entropy for the whole system in the $L/4$ : rest bipartition ($S(L/4)$) is plotted for different $\tau$ values with respect to the system size $L$. The $\tau$ values increase as the color of the curves becomes deeper. In this context, we have considered values of $\tau$ ranging from $0.06$ to $0.40$, with an increment of $0.02$ between successive $\tau$ values. We have taken $n=10$ for measuring $S$. Now, among all the $\tau$ values, i.e for $\tau=0.06,0.08,0.10,0.12,0.14,....,0,20,0.22,0.24...,0.40$, the curves are almost flat bellow $\tau=0.20$ and above $\tau=0.20$, almost linear increment of entanglement starts to arise. It is apparent that there is an area law to volume law transition around $\tau_{c}=0.20$. In panel (b) we have shown the scaling collapse (following Eq. 7) of the entanglement curves occurs when $\tau_{c}=0.2$ with exponent $\nu=1.013$.
Figure 3: Scaling collapse of height of plateau region in $R_{n}$ vs $n$ curve at $\tau_{c}=0.2$ for different system sizes at the field strength $h=1/2$. The value of $\lambda$ corresponding to height collapse is given by, $0.4317$.